Pat is 24 years older than his son Patrick. In 4 years Pat will be twice as old as Patrick. How old are they now?

Answer by ankor@dixie-net.com(22740)

Let s = son's present age
let f = father's present age
:
Pat is 24 years older than his son Patrick.
f = s + 24
:
In 4 years Pat will be twice as old as Patrick.
f + 4 = 2(s + 4)
f + 4 = 2s + 8
f = 2s + 8 - 4
f = 2s + 4
replace f with (s+24)
s + 24 = 2s + 4
24 - 4 = 2s - s
s = 20 is the son's present age
then
20 + 24 = 44 yrs father's age
:
:
Check in four years father is 48 and son is 24.